OpenStudy (anonymous):

A bag contains 6 cherry, 3 orange, and 2 lemon candies. You reach in and take 3 pieces of candy at random. The probability that you have 1 cherry candy and 2 lemon candies is

OpenStudy (anonymous):

Ans= (probability of picking cherry candy) * (probability of picking lemon candy #1) * (probability of picking lemon candy#2) = 6/11 * 2/10 * 1/9 = 2/165